OpenStudy (anonymous):

Part 1: Explain, in complete sentences, how you would use the elimination method to solve the following system of equations. (3 points) Part 2: Provide the solution to the system. (1 point) 2x + 9y = 4 3x + 7y = –7

OpenStudy (anonymous):

have a common variable for easier removal 3(2x+9y = 4) 2(3x+7y = -7) =6x+27y = 12 = 6x+14y = -14 next we eliminate variable x by subtracting the 2 equations 6x+27y = 12 -(6x+14y = -14) = 13y=26 y=2 now we found y find x by plugging value of y into any equation 2x +9(2) = 4 3x+7(2)=-7 2x+18 = 4 3x+14 = -7 2x=-14 3x= -21 x=-7 x=-7